OpenStudy (fernandez555):

Help! Find the non-extraneous solutions of sqrtx+6 - 5 = x + 1 A. x = −2 B. x = −6 C. x = −6 and x = −2 D. x = 2 I am stuck with (x+6)^1/2 = x +6

OpenStudy (anonymous):

as you have (x+6)^1/2 = x +6 squaring both sides we get x+6 = (x+6)^2 can you do it from here

OpenStudy (tkhunny):

#1 thing to do is to WRITE DOWN \(x + 6 \ge 0 \implies x \ge -6\) #2 thing to do is add 5, as you have done. #3 thing to do is think about it and stop manipulating symbols for a moment. What two numbers are their own square roots? #4 thing to do Solve x + 6 = One of them x + 6 = The other one #5 thing to do, after identifying the two possible solutions, and solving, compare them to -6.